OpenStudy (anonymous):

sketch the graph y=sec4x

OpenStudy (hba):

welcome to open study

OpenStudy (vishweshshrimali5):

@hba many new users today Its nice @samsam15 welcome to os

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3Dsec4x

OpenStudy (cherylim23):

You can write y=sec(4x) as y = 1/cos(4x) It is now a reciprocal graph of y=cos(4x) Where cos(4x) cuts the x-intercepts, there will be an asymptote, such as at x = pi/2 The graphs y=cos(4x) and y=1/cos(4x) will also intersect at the lines y = 1 or y=-1. With this, you can generally plot the graph. Watch asymptotic behaviour